OpenStudy (anonymous):

If the line passing through the points (a, 1) and (−3, 8) is parallel to the line passing through the points (5, 6) and (a + 2, 1), what is the value of a

OpenStudy (cwrw238):

8-1/-3-a = 1-6 /(a+2)-5 - 2 slopes - this equation holds if two lines are parallel

OpenStudy (anonymous):

how did you solve it?

OpenStudy (cwrw238):

solve for a: 15+ 5a = 7a + 14 - 35 2a = 36 a = 18

OpenStudy (cwrw238):

check : (18, 1) and (-3,8) slope = 8-1/-3-18 = 7/-21 = -1/3 (5,6) and (18+2, 1) = (5,6) and ( 20,-1) slope = 1-6/ 20-5 = -5/15 = -1/3 slopes are equal so lines are parallel a=18 is correct

OpenStudy (cwrw238):

do u need any more explanation?

OpenStudy (anonymous):

thank you! i was getting the problem wrong because when i solved for a i was subtracting insterad of adding i thought you had to do opposites or something